OpenStudy (anonymous):

Find the equation of the line whose slope is 2 and which passes through the point (-2, 4).

OpenStudy (anonymous):

sure. use point slope formula \[y-4=2(x+2)\]

OpenStudy (anonymous):

To do this, we need to use point slope form. This is y-y1=m(x-x1). So we plug in to get the following: y-4=2(x+2) so we get y=2(x+2)+4

OpenStudy (anonymous):

solve for y if you like \[y-4=2x+4\] \[y=2x+8\]

OpenStudy (anonymous):

point slope formula is \[y-y_1=m(x-x_1)\] and in your case \[m=2\] \[x_1=-2\] \[y_1=4\]